# Retrieving the hot CGM physics from the X-ray radial profile from eROSITA with an IlustrisTNG-based forward model

### v1

@@ 4.1 Predicting the X-ray emission from XRB
We quantify the contribution of the total $L_{X,\ \rm XRB}$ in the LS DR9 Fullphot galaxy catalogue using the model from Aird et al. (2017). Since we will later use these estimates to inform our forward models for the allowed range of point source luminosities, we adopt a TNG-independent method to predict $L_{X,\ \rm XRB}$ unbiasedly (other reasons for not using TNG also detailed in Sec. 3.1). We use uchuu, a suite of ultra-large cosmological N-body simulations (Ishiyama et al. 2021), with the galaxy catalogue from universemachine (Behroozi et al. 2019) to construct a mock for the Fullphot galaxy sample. The SFRs from universemachine are calibrated to reproduce observations. We use the half-sky lightcone, constructed in the procedure as detailed in Comparat et al. (2020), for building the mock galaxy catalogue. The mocks are generated similarly to Z24, ensuring the galaxy stellar mass function of the LS DR9 Fullphot galaxy catalogue is reproduced. Therefore, they can be reliably used for the purpose of this study. We apply Eq. 4 on the mocks to estimate the contribution of XRB in the galaxy stack, given the stellar masses and SFR of the mock galaxies. With these ingredients, we predict that the contribution from XRB alone to be $L_{X,\rm\ XRB}=(8.2_{-1.7}^{+2.2})\times 10^{39}$ ergs/s, represented by the grey hatched region in Fig. 2.
Figure 5: The fractional contribution to the total X-ray surface brightness profile of the hot CGM from central galaxies (yellow), the events around satellites probing the hot gas of their more massive host halos (green), and X-ray events from unresolved and resolved point-like sources comprising AGN and XRB (dashed grey line). The errors on the profiles are obtained from the posterior distributions of the MCMC fitting analysis
@@
X-ray emission from AGN originates around an accreting supermassive black hole (see Brandt & Alexander 2015 for a review), appearing as a point-like X-ray source with eROSITA.
We proceed to use the empirical model from Aird et al. (2013, hereafter A13) to estimate the $L_{X,\rm\ AGN}$ for a mean stellar mass $M_{\star}=10^{10.7}\ {\rm M}_{\odot}$. A13 provide a model for the probability of a galaxy hosting an AGN for a given stellar mass, $M_{\star}$, and redshift, $z$ as a function of the specific black hole accretion rate, $\lambda$ [ergs/s/${\rm M}_{\odot}$]; also called the incidence rate distribution, $\mathcal{P}(\lambda\ |\ M_{\star},z)$. The specific accretion rate, $\lambda$, of an AGN is the rate at which mass is accreting onto the supermassive black hole. The model C from A13 successfully predicts the XLF and its evolution at $1.00.2<z<$ 1.0. The specific accretion rate, $\lambda$, is related to the X-ray luminosity,
@@
where the $0.002\ M_{\star}$ factor represents the mass of the black hole, $M_{\bullet}$, and assumes correlation between $M_{\bullet}$ and the mass of the bulge, $M_{\rm bulge}$ (Marconi & Hunt 2003). Additionally, we also assume $M_{\star}\approx M_{\rm bulge}$ (A13). For the mean of our Fullphot, $M_{\star}=10^{10.7}\ {\rm M}_{\odot}$ and at the median redshift, $\langle z\rangle=0.14$, we obtain the incidence rate distribution, $\mathcal{P}(\lambda\ |\ L_{X},\ M_{\star},z)$, as a function of $L_{X,\rm\ AGN}^{2-10\ \rm keV}$ using Eq. 5. To obtain the $0.5-2$ keV mean observed X-ray luminosity, which is required to compare with the estimate from this work, we further need to convert the incidence rate distribution from $L_{X,\rm\ AGN}^{2-10\ \rm keV}$ to $L_{X,\rm\ AGN}^{0.5-2\ \rm keV}$.
@@
An additional consideration is that the optical sample used for X-ray stacking in Z24 excludes objects classified as point sources in optical. This effectively excludes the optically bright quasars, where the point-like emission strongly dominates over the host galaxy contribution. An unsolved and open question is how such optical selection criteria for AGN modify the X-ray luminosity distribution in X-rays, and addressing this is beyond the scope of this work. Nonetheless, we proceed to compute a conservative X-ray luminosity threshold to account for this exclusion of optical quasars as follows. We convert the optical r-band luminosity distribution of the Fullphot galaxy sample to the $2-10$ keV luminosity distribution using a bolometric correction factor of $2.5$ (Collin et al. 2002; Duras et al. 2020; Buchner et al. 2024). We use 10×10$\times$10 × the mean of the HXB luminosity distribution as the threshold above which the object is classified as a bright point source in the optical LS DR9 catalogue. This conservative limit excludes objects with $L_{X,\rm\ AGN}^{2-10\ \rm keV}>8\times 10^{43}$ ergs/s. We adopt this cut in the $\mathcal{P}(\lambda\ |\ L_{X},\ M_{\star},z)$ distribution as a function of the $L_{X,\rm\ AGN}^{2-10\ \rm keV}$ from A13. After applying the obscuration model from Comparat et al. (2019), we obtain $\langle L_{X,\rm\ AGN}^{0.5-2\ \rm keV}\rangle=2.7_{-2.0}^{+2.2}\times 10^{40}$ ergs/s. The sum of $L_{X,\rm\ AGN}$ computed here and $L_{X,\rm\ XRB}$ computed in the previous Sec. 4.1 is represented with the green shaded region in Fig. 2.
The large error bars on our estimate of $L_{X,\rm\ AGN}$ using the methodology discribed here are due to the uncertainties in the empirical obscuration model and the uncertainties in the incidence rate distrubution, which is poorly constrained for the low-redshift universe. The estimates here can be further improved with future works that stregthen the connection between the low luminosity X-ray AGN population with the host galaxy properties, proper knowledge mapping AGN selection functions from optical to X-ray luminosities, and better contrained obscuration models.
4.3 Using Model $3$ for interpreting the Fullphot data
In the light of the empirical estimates we obtain from Sec. 4.1 and 4.2, we compare the prediction for point source luminosities from our three forward models (based on the different halo distribution shown in Fig. 1) with the empirically allowed range of point source luminosities as shown in Fig. 2. This comparison favours model three, which has a mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$, implying that the hot CGM component allows for a point source component with a luminosity that agrees with empirical estimates from the low redshift universe. We focus our results on model three for all the following discussions of the hot CGM.
The results of fitting model $3$ to the X-ray surface brightness profile obtained by stacking on the Fullphot galaxies is shown in the left panel of Fig. 3, with the posterior distribution of the best-fit parameters shown in the right panel of Fig. 4.
4.4 X-ray emission from the hot CGM
The contribution of the hot CGM component from central galaxies is shown with the yellow line in Fig. 3 for the forward-model $3$. By integrating the area under the mean X-ray surface brightness profile from the central galaxy hot CGM component within $R_{500c}$, we obtain an X-ray luminosity, $L_{X,\ \rm CGM}=1.69_{-0.91}^{+0.28}\times 10^{40}$ ergs/s. We also show the residual plot of the per cent deviation of the data from our model, where the discrepancies are within $15\%$.
We show the fractional contribution of the various emission components in our forward model $3$ to the mean X-ray surface brightness profile upon stacking galaxies in Fig. 5. We note that at mean redshifts of $0.14$ and the underlying halo mass distribution for model $3$, the hot CGM is unresolved with an eROSITA-like PSF. Thus, at $\lesssim 40$ kpc, the hot CGM from central galaxies and the X-ray point sources emission from XRB and AGN each account for up to $40-50\%$ of the total X-ray emission budget, respectively. At larger radii $\gtrsim 40$ kpc, the contribution from the emission around satellites dominates the total X-ray emission, thereby explaining the overall flattening in the measurements. This emission around satellites arises from the more massive halos in which the satellite galaxies resides; we find the mean host halo mass of the satellite galaxies to be $M_{\rm 200m}\sim 10^{14}\ {\rm M}_{\odot}$. By integrating the area under the X-ray surface brightness profile from the emission around satellites, we obtain an X-ray luminosity, $L_{X,\ \rm SAT}=3_{-2}^{+3}\times 10^{41}$ ergs/s. We reiterate that the Fullphot galaxy catalogue is not classified into central and satellite galaxies. From our procedure of mock catalogue creation (Sec. 3.3), we predict ${S}_{\rm X,\ sat}$ contribution to the Fullphot galaxy stack. Upon fitting the forward model to the Fullphot data, we obtain $\mathcal{N}_{\rm sat}$: the renormalization parameter by which the ${S}_{\rm X,\ sat}$ contribution must be downscaled. We introduce $\mathcal{N}_{\rm sat}$ due to the overprediction of hot gas fraction in TNG in halos $\gtrsim 10^{13.5}\ {\rm M}_{\odot}$ reported in Popesso et al. (2024a) in addition to the TNG-based ${S}_{\rm X,\ sat}$ prediction being too bright (see details in Sec. 3.2). From our fitting analysis for model $3$, we find that the TNG-based ${S}_{\rm X,\ sat}$ normalization of the hot CGM must be rescaled by $0.15$ to obtain observationally consistent contributions for the Fullphot galaxy catalogue based X-ray stack.
We compare our results with the other hot CGM measurements presented in Z24, based on a different optical galaxy catalogue, namely from the SDSS spectroscopic survey. Given the spectroscopic optical information, the galaxy sample is classified into centrals and satellites (Tinker 2021), which makes it possible to empirically model the hot CGM profile from other contaminating effects (point sources and satellites). They selected $30,825$ central galaxies with spectroscopic redshifts $<0.2$ and MW-like stellar masses of $1110.5<\log(M_{\star}/{\rm M}_{\odot})<$ 11. In Z24, this SDSS-based spectroscopic sample is called the CEN sample. The resulting profile, as shown by the data points in Fig. 6, is compared with the hot CGM component (model $3$) we obtain in this work (solid yellow line). Our TNG-based forward model of the hot CGM prediction is in excellent agreement with the hot CGM measurement from Z24 at $\gtrsim 60$ kpc. At the inner radii ($\lesssim 60$ kpc), our TNG-based model $3$ overpredicts the X-ray emission. We note that the halo mass distributions of the two samples have similar mean values, where the mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$ for our forward model $3$ and the mean $M_{\rm 200m}^{\rm CEN}\sim 3\times 10^{12}\ {\rm M}_{\odot}$. However, the median $M_{\rm 200m}^{\rm CEN}\sim 1.3\times 10^{12}\ {\rm M}_{\odot}$, highlighting the spread in the halo mass distribution. This result further emphasizes the importance of the underlying halo mass distribution and the impact of the halo mass scatter introduced in stellar-mass selections when comparing hot CGM profiles across different observations and simulation-based models. For reference, we also show the forward models $1$ and $2$, which we exclude because their hot CGM component does not allow for a permissible contribution of point source luminosity in the X-ray galaxy stack (see Sec. 4.3). In addition to this shortcoming, we find that models $1$ and $2$ are discrepant with the CEN sample measurement of the hot CGM, further favouring model $3$.
@@ 4.4 X-ray emission from the hot CGM
Figure 6: Comparison of the hot gas CGM profile from our TNG-based model $3$ (solid yellow line) in this work with the hot CGM measurement from Z24 based on X-ray stacking at the optical positions of galaxies from the SDSS spectroscopic galaxy catalogue (CEN sample). The CEN sample has a mean redshift $\langle z_{\rm spec}\rangle=0.08$, which is lower than that for the Fullphot galaxy sample, $\langle z_{\rm phot}\rangle=0.14$, modelled in this work. Therefore, our models (solid yellow, dashed pink, and solid purple) are convolved with the eROSITA PSF representing $z=0.08$ to enable comparison with the CEN sample profile from Z24. Our Fullphot-based model $3$ convolved with the $\langle z_{\rm phot}\rangle=0.14$ PSF in shown by the dotted-yellow line. For reference, we also show the other two models, $1$ and $2$, with different underlying halo distributions (see text in Sec. 3.3), which we excluded in this work (see Fig. 2) as the hot CGM component did not allow for a point source component with a luminosity consistent with empirical estimates from the low redshift universe. The model $3$ from this work is in good agreement with the CEN sample hot CGM profile from Z24.

### v2

@@ 4.1 Predicting the X-ray emission from XRBs
We quantified the contribution of the total $L_{X,\ \rm XRB}$ in the LS DR9 Fullphot galaxy catalog using the model from Aird et al. (2017). Since we would later use these estimates to inform our forward models for the allowed range of point source luminosities, we adopted a TNG-independent method of predicting $L_{X,\ \rm XRB}$ unbiasedly (other reasons for not using TNG are also detailed in Sec. 3.1). We used uchuu, a suite of ultra-large cosmological N-body simulations (Ishiyama et al. 2021), with the galaxy catalog from universemachine (Behroozi et al. 2019) to construct a mock for the Fullphot galaxy sample. The SFRs from universemachine were calibrated to reproduce observations. We used the half-sky lightcone, constructed in the procedure detailed in Comparat et al. (2020), to build the mock galaxy catalog. The mocks were generated similarly to Z24, ensuring that the galaxy stellar mass function of the LS DR9 Fullphot galaxy catalog was reproduced. Therefore, they could reliably be used for the purpose of this study. We applied Eq. 4 on the mocks to estimate the contribution of XRBs in the galaxy stack, given the stellar masses and SFR of the mock galaxies. With these ingredients, we predict that the contribution from XRBs alone to be $L_{X,\rm\ XRB}=(8.2_{-1.7}^{+2.2})\times 10^{39}$ ergs/s, represented by the hatched gray region in Fig. 2777We note that the X-ray scaling relation from Lehmer et al. (2019) predicts $L_{X,\rm XRB}=(5.1_{-0.9}^{+1.6})\times 10^{39}$ erg/s for the Fullphot galaxy catalog. Nevertheless, the predictions from using the Aird et al. (2017) and Lehmer et al. (2019) scaling relations are consistent with each other, and we have verified that adopting the Lehmer relation would not change our results. .
Figure 3: Decomposition of the X-ray stack of the galaxies in the photometric sample, Fullphot, into contributions from hot gas events (centrals and satellites hosted by more massive host halos) and point sources (AGNs and XRBs). The orange data points from Z24 are described with the model from this work (shown by the solid black line). The dash-dotted orange line at $292$ kpc corresponds to the virial radius of the observational sample. The model is composed of the following: the hot CGM from central galaxies (yellow), the events around satellites probing the hot gas of their more massive host halos (green), and X-ray events from unresolved and resolved point-like sources comprising AGNs and XRBs (gray). The bottom panel shows the percentage deviation of the best-fit forward model from the data. The dash-dotted lines show the $15\%$ level.
Figure 4: Posterior probability distributions of the renormalization factor of the ${S}_{\rm X,\ sat}$ profile: $\mathcal{N}_{\rm sat}$, and the normalization of the point source component: $\mathcal{N}_{\rm ps}$, which are obtained by fitting the forward-model 33 from this work to the Fullphot data points from Z24 shown in Fig. 3. The vertical red lines in the diagonal plots correspond to the most likely value; the respective values are mentioned in the titles (refer to Tab. 1). The dashed black lines are the $68\%$ confidence interval of the marginalized distribution of the free parameters. The contour plot marks the most likely values with the red cross, and the contours correspond to the $68\%,95\%$, and $99.7\%$ confidence intervals.
@@
[…]
@@
where the $0.002\ M_{\star}$ factor represents the mass of the black hole, $M_{\bullet}$, and assumes correlation between $M_{\bullet}$ and the mass of the bulge, $M_{\rm bulge}$ (Marconi and Hunt 2003). Additionally, we also assume $M_{\star}\approx M_{\rm bulge}$ (A13). For the mean of our Fullphot, $M_{\star}=10^{10.7}\ {\rm M}_{\odot}$ and at the median redshift, $\langle z\rangle=0.14$, we obtain the incidence rate distribution, $\mathcal{P}(\lambda\ |\ L_{X},\ M_{\star},z)$, as a function of $L_{X,\rm\ AGN}^{2-10\ \rm keV}$ using Eq. 5. To obtain the 0.520.5-2 keV mean observed X-ray luminosity, which is required to compare with the estimate from this work, we further need to convert the incidence rate distribution from $L_{X,\rm\ AGN}^{2-10\ \rm keV}$ to $L_{X,\rm\ AGN}^{0.5-2\ \rm keV}$.
An important factor that comes into play when performing a conversion from 2102-10 keV (hard X-ray band; HXB) to the 0.520.5-2 keV (soft X-ray band; SXB) luminosity is the intrinsic obscuration of the AGN. Our estimate of $L_{X,\ \rm AGN}$ represents the contribution from the obscured AGNs and the observed unobscured Type 1 AGNs, the dominating component at the luminosity range under concern (see e.g., Hasinger 2008). We used the Comparat et al. (2019) empirical obscuration model to obtain the observed HXB-to-SXB luminosity conversion; they self-consistently build an obscuration model based on observation works (Ricci et al. 2017; Buchner and Bauer 2017; Ueda et al. 2014; Aird et al. 2015; Buchner et al. 2015). The Comparat et al. (2019) model is implemented on the uchuu simulations (introduced in Sec. 4.1), and we obtain the HXB to SXB conversion as a function of $L_{X,\rm\ AGN}^{2-10\ \rm keV}$. Finally, we obtain the desired A13-based $\mathcal{P}(\lambda\ |\ L_{X},\ M_{\star},z)$ distribution as a function of the $L_{X,\rm\ AGN}^{0.5-2\ \rm keV}$. The expectation value is the obtained as follows: $\langle L_{X,\rm\ AGN}^{0.5-2\ \rm keV}\rangle=\int\mathcal{P}(\lambda\ |\ L_{X},\ M_{\star},z)L_{X,\rm\ AGN}^{0.5-2\ \rm keV}d\lambda$.
An additional consideration is that the optical sample used for X-ray stacking in Z24 excludes objects classified as point sources in optical. This effectively excludes the optically bright quasars, where the point-like emission strongly dominates over the host galaxy contribution. An unsolved and open question is how such optical selection criteria for AGNs modify the X-ray luminosity distribution in X-rays, and addressing this is beyond the scope of this work. Nonetheless, we proceeded to compute a conservative X-ray luminosity threshold to account for this exclusion of optical quasars as follows. We converted the optical r-band luminosity distribution of the Fullphot galaxy sample to the 2102-10 keV luminosity distribution using a bolometric correction factor of $2.5$ (Collin et al. 2002; Duras et al. 2020; Buchner et al. 2024). We used $10\times$ the mean of the HXB luminosity distribution as the threshold above which the object is classified as a bright point source in the optical LS DR9 catalog. This conservative limit excludes objects with $L_{X,\rm\ AGN}^{2-10\ \rm keV}>8\times 10^{43}$ ergs/s. We adopted this cut in the $\mathcal{P}(\lambda\ |\ L_{X},\ M_{\star},z)$ distribution as a function of the $L_{X,\rm\ AGN}^{2-10\ \rm keV}$ from A13. After applying the obscuration model from Comparat et al. (2019), we obtained $\langle L_{X,\rm\ AGN}^{0.5-2\ \rm keV}\rangle=2.7_{-2.0}^{+2.2}\times 10^{40}$ ergs/s. The sum of $L_{X,\rm\ AGN}$ computed here and $L_{X,\rm\ XRB}$ computed in the previous Sec. 4.1 is represented with the shaded green region in Fig. 2.
The large error bars on our estimate of $L_{X,\rm\ AGN}$ using the methodology described here are due to the uncertainties in the empirical obscuration model and the uncertainties in the incidence rate distribution, which is poorly constrained for the low-redshift Universe. The estimates here can be further improved with future works that strengthen the connection between the low-luminosity X-ray AGN population and the host galaxy properties, proper knowledge mapping AGN selection functions from optical to X-ray luminosities, and better constrained obscuration models.
Figure 5: Fractional contribution of centrals, satellites, and point sources to the total X-ray surface brightness profile of the hot CGM. The yellow line shows the contribution from central galaxies, the green line shows the events around satellites probing the hot gas of their more massive host halos, and the dashed gray line shows the X-ray events from unresolved and resolved point-like sources comprising AGNs and XRBs. The errors on the profiles were obtained from the posterior distributions of the MCMC fitting analysis.
4.3 Using Model $3$ to interpret the Fullphot data
In the light of the empirical estimates we obtain from Sec. 4.1 and 4.2, we compare the prediction for point source luminosities from our three forward models (based on the different halo distribution shown in Fig. 1) with the empirically allowed range of point source luminosities as shown in Fig. 2. This comparison favors model three, which has a mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$, implying that the hot CGM component allows for a point source component with a luminosity that agrees with empirical estimates from the low-redshift Universe. We focus our results on model three for all the following discussions of the hot CGM. The results of fitting model $3$ to the X-ray surface brightness profile obtained by stacking on the Fullphot galaxies are shown in Fig. 3, with the posterior distribution of the best-fit parameters shown in Fig. 4.
4.4 X-ray emission from the hot CGM
The contribution of the hot CGM component from central galaxies is shown with the yellow line in Fig. 3 for the forward-model $3$. By integrating the area under the mean X-ray surface brightness profile from the central galaxy hot CGM component within $R_{500c}$, we obtain an X-ray luminosity, $L_{X,\ \rm CGM}=1.69_{-0.91}^{+0.28}\times 10^{40}$ ergs/s. We also show the residual plot of the percentage deviation of the data from our model, where the discrepancies are within $15\%$.
We show the fractional contribution of the various emission components in our forward model $3$ to the mean X-ray surface brightness profile upon stacking galaxies in Fig. 5. We note that at mean redshifts of $0.14$ and the underlying halo mass distribution for model $3$, the hot CGM is unresolved with an eROSITA-like PSF. Thus, at $\lesssim 40$ kpc, the hot CGM from central galaxies and the X-ray point sources emission from XRBs and AGNs each account for up to $40-50\%$ of the total X-ray emission budget, respectively.
We compare our results with the other hot CGM measurements presented in Z24, based on a different optical galaxy catalog, namely from the SDSS spectroscopic survey. Given the spectroscopic optical information, the galaxy sample is classified into centrals and satellites (Tinker 2021), which makes it possible to empirically model the hot CGM profile from other contaminating effects (point sources and satellites). They selected $30,825$ central galaxies with spectroscopic redshifts <0.2<0.2 and MW-like stellar masses of $10.5<\log(M_{\star}/{\rm M}_{\odot})<11$. In Z24, this SDSS-based spectroscopic sample is called the CEN sample. The resulting profile, as shown by the data points in Fig. 6, is compared with the hot CGM component (model $3$) we obtain in this work (solid yellow line). Our TNG-based forward model of the hot CGM prediction is in excellent agreement with the hot CGM measurement from Z24 at $\gtrsim 60$ kpc. At the inner radii ($\lesssim 60$ kpc), our TNG-based model $3$ overpredicts the X-ray emission. We note that the halo mass distributions of the two samples have similar mean values, where the mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$ for our forward model $3$ and the mean $M_{\rm 200m}^{\rm CEN}\sim 3\times 10^{12}\ {\rm M}_{\odot}$. However, the median $M_{\rm 200m}^{\rm CEN}\sim 1.3\times 10^{12}\ {\rm M}_{\odot}$, highlighting the spread in the halo mass distribution. This result further emphasizes the importance of the underlying halo mass distribution and the impact of the halo mass scatter introduced in stellar-mass selections when comparing hot CGM profiles across different observations and simulation-based models. For reference, we also show the forward models $1$ and $2$, which we exclude because their hot CGM component does not allow for a permissible contribution of point source luminosity in the X-ray galaxy stack (see Sec. 4.3). In addition to this shortcoming, we find that models $1$ and $2$ are discrepant with the CEN sample measurement of the hot CGM, further favoring model $3$.
4.5 X-ray emission from satellite’s host halos
@@
Upon fitting our forward model to the Fullphot stack, we introduced a renormalization parameter, $\mathcal{N}_{\rm sat}$, which rescales the ${S}_{\rm X,\ sat}$ contribution to match the observations. This scaling was necessary due to the known overprediction of hot gas fractions in TNG300 for halos above $10^{13.5}\ {\rm M}_{\odot}$ (Popesso et al. 2024a), leading to an overly bright ${S}_{\rm X,\ sat}$ component (see Sec. 3.2). Importantly, the shape of this satellite-related emission component is independent of the satellite fraction, as demonstrated in Fig. 9 in Shreeram et al. (2025). From our fitting analysis for model 3, we find that the TNG-based fiducial ${S}_{\rm X,\ sat}$ normalization of the hot CGM must be rescaled by $0.15$ to provide an observationally consistent contribution for the Fullphot galaxy catalog based X-ray stack. We highlight that this renormalization does not affect the hot CGM contribution from centrals, which remains unchanged. Therefore, $\mathcal{N}_{\rm sat}$ effectively provides a quantitative correction factor, illustrating how the overabundance of hot gas in massive halos biases the observable MW-mass stacked X-ray profiles. We conclude that the inclusion and forward-modeling of the satellite’s host halo component is required to physically interpret the Fullphot data.
@@
Figure 6: Comparison of the hot gas CGM profile from our TNG-based model 33 (solid yellow line) in this work with the hot CGM measurement from Z24 based on X-ray stacking at the optical positions of galaxies from the SDSS spectroscopic galaxy catalog (CEN sample). The CEN sample has a mean redshift $\langle z_{\rm spec}\rangle=0.08$, which is lower than that for the Fullphot galaxy sample, $\langle z_{\rm phot}\rangle=0.14$, modeled in this work. Therefore, our models (solid yellow, dashed pink, and solid purple) are convolved with the eROSITA PSF representing $z=0.08$ to enable comparison with the CEN sample profile from Z24. Our Fullphot-based model 33 convolved with the $\langle z_{\rm phot}\rangle=0.14$ PSF in shown by the dotted yellow line. For reference, we also show the other two models, 11 and $2$, with different underlying halo distributions (see text in Sec. 3.3), which we excluded in this work (see Fig. 2) as the hot CGM component did not allow for a point source component with a luminosity consistent with empirical estimates from the low-redshift Universe. The model $3$ from this work is in good agreement with the CEN sample hot CGM profile from Z24.